Circular No. 6487

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 960720 J. Greiner, Max-Planck-Institut (MPI) für Extraterrestrische Physik,

Garching; H.-J. Hagen, Hamburg Observatory; and A. Heines, MPI group at Jena University, report: “A 6300-s ROSAT HRI observation on Aug. 31 reveals the existence of eight x-ray sources above the 3σ level within the 10′- radius error box of GRB 960720 (IAUC 6467). All eight x-ray sources have intensities below the ROSAT all-sky-survey sensitivity threshold and their brightness history thus remains unknown. Further optical and x-ray observations are necessary to search for variability patterns that might suggest a relationship of one of these sources to GRB 960720. Three of these eight x-ray sources are located within the 1′-radius error box of the x-ray source detected with BeppoSAX (IAUC 6480) and ASCA (IAUC 6481), 2′.9 off the GRB-error-box center. Each of these three x-ray-source error boxes (± 10′′) contains only one optical object down to the POSS limit (a finding chart can

be found at http://www.rosat.mpe-garching.mpg.de/~jcg/grb.html).

Optical spectra in the range 350–900 nm, obtained on Oct. 3 and 4 with the MPI für Astronomie 2.2-m telescope at Calar Alto, show strong and broad emission lines at 373.8 and 546.1 nm in RX J1729.5+4904 and at 356.4 and 525.5 nm in RX J1729.3+4904. These two lines in each object are interpreted as C III] at 190.9 nm and Mg II at 279.8 nm, resulting in redshifts of 0.96 and 0.87 for these objects. The optical intensities (magnitudes B = 20.8 and 20.0) and redshifts suggest a Seyfert-galaxy nature for both objects. The optical object seen on the POSS O plate within the error circle of the third x-ray source (RX J1729.4+4905) is currently fainter than B = 22. Assuming a power-law model for the x-ray spectrum with photon index −1.9 and only galactic absorption (2.5 × 10^{20} cm^{−2}), the HRI countrate of 0.0017 count/s for RX J1729.5+4904 transforms into an unabsorbed x-ray flux of 9.4 × 10^{−14} erg cm^{−2} s^{−1} in the range 0.1–2.4 keV. The flux of the single source seen with BeppoSAX and ASCA is consistent with the sum of all three HRI x-ray sources (0.0011 count/s for RX J1729.3+4904 and 0.0008 count/s for RX J1729.4+4905).”

COMET C/1996 N1 (BREWINGTON) Visual m_{1} estimates by K. Hornoch, Lelekovice, Czech Republic (B =

10×80 binoculars; L = 0.35-m reflector): Sept. 2.81 UT, 8.4 (B); 11.85, 9.0 (B); 19.83, 9.3 (B); Oct. 3.83, 10.1 (L); 7.77, 11.0 (L).

1996 October 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green